1) A fish can swim at 14m/s with the current and at 6m/s against it. Find the speed of the current and the speed of the fish in still water.
2) If the numerator and denominator of a fraction are both decreased by one the fraction becomes 2/3. If the numerator and denominator are both increased by one the fraction becomes 3/4. Find the original fraction.

Answer by mananth(16949)

let speed of fish in still water be x
let speed of current be y m/s
..
x+y =14
x-y =6
add the equations
2x=20
/2
x= 10 m/s
y = 4m/s
...
m.ananth@hotmail.ca
...
let the numerator be x
denominator be y
...
(x-1)/(y-1) =2/3
3(x-1)=2(y-1)
3x-3=2y-2
3x-2y =1.................1
...
(x+1)/(y+1)=3/4
4(x+1)=3(y+1)
4x+4 = 3y+3
4x-3y=-1................2
multiply equation 1 by -3 & equation (2) by 2
-9x+6y=-3
8x-6y=-2
add i
-x=-5
x=5
..
plug value of x in (1)
3*(5)-2y = 1
-2y= -14
/-2
y=7
..
Fraction = 5/7
...
m.ananth@hotmail.ca